# BPINN-EM-Post: Stochastic Electromigration Damage Analysis in the Post-Void Phase based on Bayesian Physics-Informed Neural Network

### v1

@@ Abstract
In contrast to the assumptions of most existing Electromigration (EM) analysis tools, the evolution of EM-induced stress is inherently non-deterministic, influenced by factors such as input current fluctuations and manufacturing non-idealities. Traditional approaches for estimating stress variations typically involve computationally expensive and inefficient Monte Carlo simulations with industrial solvers, which quantify variations using mean and variance metrics. In this work, we introduce a novel machine learning-based framework, termed BPINN-EM-Post, for efficient stochastic analysis of EM-induced post-voiding aging processes. This new approach integrates closed-form analytical solutions with a Bayesian Physics-Informed Neural Network (BPINN) framework to accelerate the analysis for the first time. The closed-form solutions enforce physical laws at the individual wire segment level, while the BPINN ensures that physics constraints at inter-segment junctions are satisfied and stochastic behaviors are accurately modeled. By reducing the number of variables in the loss functions through the use of analytical solutions, our method significantly improves training efficiency without accuracy loss and naturally incorporates variational effects. Additionally, the analytical solutions effectively address the challenge of incorporating initial stress distributions in interconnect structures during post-void stress calculations. Numerical results demonstrate that BPINN-EM-Post achieves over 240×240$\times$240 × speedup compared to Monte Carlo simulations using the FEM-based COMSOL solver and more than 65×65$\times$65 × speedup compared to Monte Carlo simulations using the FDM-based EMSpice method.
Index Terms:
Post-void Electromigration (EM), Physics informed neural network (PINN), Bayesian networks
@@ I Introduction
Electromigration (EM) is a phenomenon where metal atoms within interconnects migrate as a result of interactions with high-current-carrying electrons. This atomic migration generates compressive stress at the anode and tensile stress at the cathode. When the stress exceeds a critical threshold, void formation near the cathode or hillock formation near the anode may occur, potentially causing circuit failure. The EM lifetime of interconnect wires is commonly characterized by the mean time to failure (MTTF), reflecting the inherent variability of the process [1].
@@
Traditionally, the mean time to failure (MTTF) or reliability of an interconnect wire is linked to the wire’s current density through well-established Electromigration (EM) models based on the work of Black and Blech [2, 3]. However, these single-wire segment EM models have been increasingly criticized for being overly conservative, as they treat each wire segment in isolation. Recent research demonstrates that stress evolution among wire segments within an EM tree—confined by the metal atom liner and barriers of the interconnect—is highly correlated and should be analyzed collectively [4, 5]. As a result, numerous physics-based analytical and numerical solutions have been proposed to address the limitations of Black’s model [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. These methods primarily focus on solving Korhonen’s partial differential equation (PDE) to model EM-induced stress evolution [22]. Nevertheless, traditional numerical techniques for solving Korhonen’s equation and related PDEs remain computationally challenging. Furthermore, stochastic physics-based EM analyses using Monte Carlo simulations are prohibitively expensive.
@@
Recently, a machine learning-based method known as physics-informed neural networks (PINN) has been developed to facilitate the learning and integration of physical laws represented by nonlinear partial differential equations (PDEs) in complex physical, biological, and engineering systems [23, 24]. The PINN approach has been utilized to solve Korhonen’s equation, showing promising results for Electromigration (EM) stress modeling [25, 26, 27]. In this framework, the physical laws, boundary conditions, and initial conditions of the PDEs are explicitly enforced through loss functions within the neural networks. This method has demonstrated effectiveness for small-scale PDE problems with a limited number of variables. Recently a hierarchical PINN scheme has been introduced to accelerate EM analysis [26, 28]. This hierarchical approach addresses stress evolution in multi-segment interconnects in two stages. First, parameterized surrogate models for stress evolution in individual wire segments are constructed using either neural networks or analytical solutions. In the second stage, the PINN scheme is employed to enforce stress continuity and atomic flux conservation across all segments of a multi-segment interconnect wire. This strategy significantly reduces the number of variables in the second stage of PINN training, thereby enhancing computational efficiency. However, these PINN-based methods have yet to address the variability in EM effects.
@@
Bayesian Neural Networks (BNNs) have emerged as an important class of deep neural networks due to their robust capabilities in uncertainty quantification, handling limited datasets, and preventing overfitting [29]. Recently, Lamichhane et al. introduced a hierarchical approach that integrates a pre-trained Bayesian network with a physics-informed neural network (PINN) framework to estimate Electromigration (EM) stress variations, with a particular focus on the nucleation phase [21]. Extending BNNs to analyze the post-voiding phase presents significant challenges, as it requires incorporating the initial stress distribution, a task that is highly complex and often infeasible for even simple neural networks [28]. This work addresses this challenge by leveraging analytical solutions for stress in single wire segments, ensuring accuracy through the enforcement of stress continuity and atomic flux conservation within the Bayesian PINN (BPINN) framework. Furthermore, the Bayesian component in BPINN facilitates a variational analysis of EM stress solutions. Our new and significant contributions of this work are as follows:
@@
  • We propose a novel hierarchical PINN framework, termed BPINN-EM-Post, to address the stochastic stress evolution problem for Electromigration (EM) damage assessment in multi-segment interconnect wires during the post-voiding phase. This framework, developed for the first time, integrates analytical solutions with a Bayesian Physics-Informed Neural Network (BPINN) for efficient training and accurate variational modeling.
@@
  • Analytical solutions are utilized to precisely estimate the EM stress distribution within individual segments, whether they contain voids or not. Unlike previous methods that relied on machine modeling, our approach uses analytic solutions to overcome the complexities associated with incorporating all possible initial stress distributions from the nucleation phase into the post-voiding phase. This approach also reduces the number of variables within the hierarchical PINN framework, enhancing computational efficiency.
@@
  • The BPINN framework is specifically applied at the boundaries of single wire segments, working in conjunction with the analytical solutions. By using fewer variables in the loss function, the BPINN framework enforces key physical conditions, such as stress continuity and atomic flux conservation, while optimizing the atomic flux. This significantly accelerates the posterior sampling process, reducing the training time and enhancing overall efficiency.
@@
  • Our numerical results show that BPINN-EM-Post achieves a computational speedup of over 240×240$\times$240 × compared to Monte Carlo simulations using FEM-based COMSOL, and more than 65×65$\times$65 × compared to Monte Carlo simulations using the FDM-based EMSpice method.
The paper is organized as follows: Section II reviews existing works on physics-based EM model and its analysis techniques and recently proposed stochastic EM analysis methods. The theoretical concepts used in our method are explained in Section III. Section IV proposes the BPINN-EM-Post with detailed description of the two-step PINN scheme considering the variational modeling using Bayesian networks. Experimental results are presented in Section V. Finally, section VI concludes this paper.
II Review of relevant works
The impact of variational or stochastic electromigration (EM) on the lifespan of wires has been explored in previous studies. Research in [30] highlights that as technology advances, both the mean time to failure and the distribution of failure times deteriorate, posing ongoing challenges for advanced interconnects. Early work in [1] introduced a method that incorporates variability into EM analysis. This approach, however, examines a single wire segment in isolation and uses both analytical and numerical techniques to compute the time to failure. It considers two forms of variability: circuit-level variability, where the resistance of wires is treated as a random variable (RV), and wire geometry-level variability, where imperfections such as bumps and necking in wire geometries are RVs due to lithography-induced variations. The research in [31] looks at both global and local process variations in power grids, applying a Hermite polynomial chaos method for variational EM analysis and final lifetime estimation using Black’s EM model. In [32], a comprehensive analysis of the entire power grid is conducted, taking into account the inherent redundancy of grids by solving Korhonen’s PDE using finite difference techniques. Here, EM diffusivity is treated as an RV, and a Monte Carlo approach is employed for the variational analysis, supplemented by some ad-hoc acceleration techniques. The study in [33] primarily focuses on variability in input currents, where stochastic current models are used to model currents in functional blocks, and variance or covariance calculations are done via matrix-solving for entire power grids, which remains computationally expensive.
In recent work, Yang et al. introduced a Bayesian physics-informed neural network (PINN) framework aimed at quantifying noise in observations and addressing the overfitting challenges typical of standard PINNs [34]. They demonstrate the effectiveness of their approach on standard, relatively straightforward problems. However, their method applies Bayesian-based sampling to the entire PINN during training, which can be quite time-consuming, especially for large-scale domains such as multi-segment interconnect structures.
Lamichhane et al proposed a hierarchical approach to estimate the variations in EM stress in [21]. In this work they hierarchically combined a pretrained Bayesian network and PINN framework. This approach however is only limited to the EM stress in nucleation phase. For the stress evolution in post-voiding phase, we need to incorporate the initial stress distribution. The intial stress distribution is too complex to be mapped using a simple neural network [28]. In this work we solve this challenge by utilizing analytical solution to get stress results on single wire segments. The accuracy of stress results is ensured by the enforcement of stress continuty and atomic flux conservation using BPINN framework. More importantly, the Bayesian Network used in BPINN framework facilitates the variational analysis on the EM stress solutions.
III Background
III-A EM stress evolution in postvoiding phase
In confined metal wires subjected to high current densities, electromigration occurs as atoms are migrated from the cathode to the anode due to interactions between electrons and metal atoms [2]. Over time, this atom migration can lead to a scarcity of atoms at the cathode and an excess at the anode, potentially forming voids and hillocks that compromise interconnect functionality.
@@ III-A EM stress evolution in postvoiding phase
Various methods have been developed to evaluate the electromigration-induced reliability of metal interconnects. Among these, Blech’s limit [3] and Black’s Mean Time To Failure (MTTF) equation [2] are traditional approaches for reliability assessment. Despite their utility, these methods are becoming less popular as Black’s MTTF is limited to individual line segments and neither method accounts for the evolution of hydrostatic stress. Korhonen’s equation [35] addresses these limitations by describing electromigration in confined metal wires through a set of partial differential equations (PDEs), and extends to model hydrostatic stress in multi-segment interconnects [36, 17]. The equation for the nucleation phase in multi-segment interconnects is given by:
$\frac{\partial\sigma_{ij}(x,t)}{\partial t}=\frac{\partial}{\partial x}\left[\kappa_{ij}(\frac{\partial\sigma_{ij}(x,t)}{\partial x}+G_{ij})\right],t>0$ (1)
@@
Here, $\sigma_{ij}(x,t)$ represents the stress in the interconnect s egment $ij$ connecting nodes $i$ and $j$. In equation  (1), $G_{ij}$ denotes the electromigration (EM) driving force in segment $ij$, calculated as $G_{ij}=\frac{e\rho J_{ij}Z^{*}}{\Omega}$, where $J_{ij}$ is the current density in segment $ij$. The stress diffusivity $\kappa_{ij}$ is defined by $\kappa_{ij}=D_{a}B\Omega/(k_{B}T)$, where $D_{a}$ is the effective atomic diffusion coefficient, $B$ is the effective bulk modulus, $k_{B}$ is Boltzmann’s constant, $T$ is the absolute temperature, and $E_{a}$ is the EM activation energy. Here, $e$ is the electron charge, $\rho$ is resistivity, and $Z^{*}$ is the effective charge.
The first boundary condition (BC) in  (1) enforces stress continuity at inter-segment junctions, specifically at $x=x_{r}$. The second BC addresses atomic flux conservation at these junctions, while the third BC applies to blocking terminal boundaries $x=x_{b}$, ensuring zero atomic flux. The inward unit normal direction at the interior junction node $x_{r}$ on segment $ij$ is represented by $n_{r}$. The initial condition (IC) specifies that the initial stress distribution in segment $ij$ is given by $\sigma_{ij,T}$.
In multisegment interconnect trees, void nucleation occurs at the cathode node when the steady-state nucleation stress surpasses the critical stress $\sigma_{crit}$. The moment this happens marks the nucleation time, denoted as $t_{nuc}$. Beyond this time, $t>t_{nuc}$, the void enlarges and the dynamics of stress evolution diverge from those observed during the nucleation phase. Once nucleated, the stress at the interface of the void dramatically drops to zero due to a significant stress gradient [37]. For the segment of the interconnect where the void originates, the equations applicable in the post-voiding phase are outlined in (LABEL:eq:korhonen_equation_single_wire_postvoid[37]. This equation presumes the $x=0$ node as the cathode node where the void forms. Should a void nucleate at $x=L$, the boundary conditions (BCs) are adjusted to reflect this scenario.
@@
$BC:\frac{\partial\sigma(0,t)}{\partial x}=\frac{\sigma(0,t)}{\delta},x=0,t>0$
$BC:\frac{\partial\sigma(0,t)}{\partial x}=-G,x=L,t>0$
@@
The equation described posits that the void interface possesses an effective thickness, $\delta$, which is negligible relative to other dimensions within the interconnect tree. The examination of stress following void formation counts time, $t$, starting from the nucleation time, $t_{nuc}$. It is assumed that the initial stress distribution at this stage corresponds to the stress levels at the moment of void nucleation. This methodology facilitates the simulation of stress evolution within a single confined wire subsequent to the occurrence of void nucleation.
In multisegment interconnect trees, it is possible for multiple cathode nodes to surpass the critical stress threshold. In such instances, a void forms at the node experiencing the highest stress [38]. For analyzing multisegment interconnects during the post-voiding phase, equation (LABEL:eq:korhonen_equation_single_wire_postvoid) is applied to the segment that contains the void. Segments without voids are addressed using equation (1). To accurately determine the stress profile across the entire interconnect tree, boundary conditions (BC) that ensure stress continuity and atomic flux conservation must be implemented  [38].
III-B Estimation of variations in EM stress
Typically, we solve combinations of equations (1) and  (LABEL:eq:korhonen_equation_single_wire_postvoid) to determine the post-voiding phase transient stress distribution across general multisegment interconnects. By comparing the results with the critical stress, we then predict the time of failure based on the void nucleation time $t_{nuc}$. This conventional deterministic approach, which averages the DC input current, does not account for variations in electromigration (EM) stress caused by fluctuations in input current, manufacturing processes, temperature, and diffusivity. Consequently, this method tends to yield overly optimistic and unrealistic predictions of component lifetimes [33].
If we define $\mathbf{x}$ as the vector of electrical and physical parameters influencing EM stress development, and $v$ as the error from these variations, the uncertainty and noise in these parameters can be treated as random variables within their respective spaces, i.e., $\mathbf{x}\in\mathbb{X}$ and $v\in\mathbb{V}$. Based on the works in  [33], it is posited that the solutions for EM stress are Gaussian distributed, centered around the true, but obscured, values:

### v2

# BPINN-EM-Post: Bayesian Physics-Informed Neural Network based Stochastic Electromigration Damage Analysis in the Post-void Phase

@@ Abstract
In contrast to the assumptions of most existing Electromigration (EM) analysis tools, the evolution of EM-induced stress is inherently non-deterministic, influenced by factors such as input current fluctuations and manufacturing non-idealities. Traditional approaches for estimating stress variations typically involve computationally expensive and inefficient Monte Carlo simulations with industrial solvers, which quantify variations using mean and variance metrics. In this work, we introduce a novel machine learning-based framework, termed BPINN-EM-Post, for efficient stochastic analysis of EM-induced post-voiding aging processes. For the first time, our new approach integrates closed-form analytical solutions with a Bayesian Physics-Informed Neural Network (BPINN) framework to accelerate the analysis. The closed-form solutions enforce physical laws at the individual wire segment level, while the BPINN ensures that physics constraints at inter-segment junctions are satisfied and stochastic behaviors are accurately modeled. By reducing the number of variables in the loss functions through utilizing analytical solutions, our method significantly improves training efficiency without accuracy loss and naturally incorporates variational effects. Additionally, the analytical solutions effectively address the challenge of incorporating initial stress distributions in interconnect structures during post-void stress calculations. Numerical results demonstrate that BPINN-EM-Post achieves over $240\times$ and more than $67\times$ speedup compared to Monte Carlo simulations using the FEM-based COMSOL solver and FDM-based EMSpice, respectively, with marginal accuracy loss.
@@ I Introduction
Electromigration (EM) is a phenomenon where metal atoms within interconnects migrate due to interactions with high-current-carrying electrons, and considered as a major concern to VLSI interconnect. This atomic migration generates compressive stress at the anode and tensile stress at the cathode. When the stress exceeds a critical threshold, void formation near the cathode or hillock formation near the anode may occur, potentially causing circuit failure. The EM lifetime of interconnect wires is commonly characterized by the mean time to failure (MTTF), reflecting the inherent variability of the process [1].
@@
Traditionally, the MTTF or reliability of an interconnect wire is linked to the wire’s current density through well-established EM models by Black and Blech [2, 3]. However, these single-wire segment EM models have been increasingly criticized for being overly conservative, as they treat each wire segment in isolation. Recent research demonstrates that stress evolution among wire segments within an EM tree, is highly correlated and should be analyzed collectively [4, 5]. Numerous physics-based analytical and numerical solutions have been proposed to address the above-mentioned limitations [6, 7, 8, 9, 10, 11, 12, 13, 14]. These methods primarily focus on solving Korhonen’s partial differential equation (PDE) to model EM-induced stress evolution [15]. Nevertheless, traditional numerical techniques for solving Korhonen’s equation and related PDEs remain computationally costly. Furthermore, stochastic physics-based EM analyses using Monte Carlo simulations are prohibitively expensive for numerical solvers such as COMSOL [16] and EMSpice [13], which demand multiple samplings to estimate variation distributions.
@@
In recent years, a machine learning-based method known as physics-informed neural networks (PINN) has been developed to facilitate the learning and integration of physical laws represented by nonlinear PDEs in complex physical, biological, and engineering systems [17, 18]. PINN approach has been utilized to solve Korhonen’s equation, showing promising results for EM stress modeling [19, 20, 21]. In this framework, physical laws, boundary conditions (BCs), and initial conditions (ICs) of the PDEs are explicitly enforced through loss functions within the neural networks. This approach has demonstrated effectiveness on small-scale PDE problems. Recently, hierarchical PINN schemes have been introduced to accelerate EM analysis [20, 22]. These approaches divide the stress evolution problem in multi-segment interconnects into two stages, thereby significantly reducing the number of variables, as well as enhancing computational efficiency and scalability to large-scale interconnect structures. However, these PINN-based methods have yet to tackle the variability in EM effects.
@@
Bayesian Neural Networks (BNNs) have emerged as an important class of neural networks due to their robust capabilities in uncertainty quantification, handling limited datasets, and preventing overfitting [23]. Lamichhane et al. introduced a hierarchical approach that integrates a pre-trained Bayesian network with a PINN framework to estimate EM stress variations, with a particular focus on the nucleation phase [24]. However, extending BNNs to analyze the post-voiding phase poses significant challenges, as it requires incorporating the initial stress distribution, a task that is highly complex and often infeasible for even simple neural networks [22].
In this work, we address the above-mentioned challenges by leveraging analytical solutions for stress in single wire segments, ensuring accuracy through the enforcement of stress continuity and atomic flux conservation within the Bayesian PINN (BPINN) framework. Furthermore, the Bayesian component in BPINN facilitates a variational analysis of EM stress solutions. Our major contributions are as follows:
@@
  • We propose a novel hierarchical PINN framework termed BPINN-EM-Post to address the stochastic stress evolution problem for EM damage assessment in multi-segment interconnect wires during the post-voiding phase. This framework, developed for the first time, integrates analytical solutions with a BPINN for efficient training and accurate variational modeling.
@@
  • Analytical solutions are utilized to precisely estimate the EM stress distribution within individual segments, regardless of void formulation. Unlike prior arts that rely on machine learning modeling, analytical approach overcomes the complexities associated with incorporating all possible initial stress distributions from nucleation phase into the post-voiding phase, enhances computational efficiency by reducing the number of variables within the framework.
@@
  • The BPINN framework is specifically applied at the boundaries of single wire segments, working in conjunction with the analytical solutions. By using fewer variables in the loss function, the BPINN framework enforces key physical conditions, such as stress continuity and atomic flux conservation, while optimizing the atomic flux, which significantly accelerates the posterior sampling process.
@@
  • Numerical results show that our proposed BPINN-EM-Post achieves a computational speedup of over $240\times$ compared to Monte Carlo simulations using FEM-based COMSOL, and more than $67\times$ acceleration compared to Monte Carlo simulations using the FDM-based EMSpice method, with only negligible (less than $\sim 1\%$) error rate.
II Preliminaries
II-A EM Stress Evolution in Post-voiding Phase
@@
In confined metal wires subjected to high current densities, EM occurs as atoms are migrated from the cathode to the anode due to interactions between electrons and metal atoms [2]. Over time, this atom migration may potentially cause void and hillock formulations that compromise interconnect functionality. Various methods have been developed to evaluate the EM-induced reliability of metal interconnects. Among these, Blech’s limit [3] and Black’s MTTF [2] are traditional approaches for reliability assessment. Despite their utility, these methods are becoming less popular as Black’s MTTF is limited to individual line segments and neither method accounts for the evolution of hydrostatic stress. Korhonen’s equation [15] addresses these limitations by describing EM in confined metal wires through a set of PDEs, and extends to model hydrostatic stress in multi-segment interconnects [25, 12]. The equation for the nucleation phase in multi-segment interconnects is given by Eq. (1), where $\sigma_{ij}(x,t)$ represents the stress in the interconnect segment $ij$ connecting nodes $i$ and $j$. In Eq. (1), $G_{ij}$ denotes the EM driving force in segment $ij$, calculated as $G_{ij}=\frac{e\rho J_{ij}Z^{*}}{\Omega}$, where $J_{ij}$ is the current density in segment $ij$.
$PDE:$ $\frac{\partial\sigma_{ij}(x,t)}{\partial t}=\frac{\partial}{\partial x}\left[\kappa_{ij}(\frac{\partial\sigma_{ij}(x,t)}{\partial x}+G_{ij})\right],\ t>0$ (1)
@@
The stress diffusivity denoted as $\kappa_{ij}$ is defined by $\kappa_{ij}=D_{a}B\Omega/(k_{B}T)$, where $D_{a}$ is the effective atomic diffusion coefficient, $B$ is the effective bulk modulus, $k_{B}$ is Boltzmann’s constant, $T$ is the absolute temperature, and $E_{a}$ is the EM activation energy. $e$ is the electron charge, $\rho$ is resistivity, and $Z^{*}$ is the effective charge.
The first BC in Eq. (1) enforces stress continuity at inter-segment junctions, specifically at $x=x_{r}$. The second BC addresses atomic flux conservation at these junctions, while the third BC applies to blocking terminal boundaries $x=x_{b}$, ensuring zero atomic flux. The inward unit normal direction at the interior junction node $x_{r}$ on segment $ij$ is represented by $n_{r}$. The IC specifies that the initial stress distribution in segment $ij$ is given by $\sigma_{ij,T}$.
In multi-segment interconnect trees, void nucleation occurs at the cathode when the steady-state nucleation stress surpasses the critical stress $\sigma_{crit}$. This moment marks the nucleation time, denoted as $t_{nuc}$. Beyond this time, $t>t_{nuc}$, the void enlarges and the dynamics of stress evolution diverge from nucleation phase. Once nucleated, the stress at the interface of the void dramatically drops to zero [26]. For the segment where the void originates, the equations applicable in the post-voiding phase are outlined in Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid[7], which presumes the $x=0$ node as the cathode where the void forms. Should a void nucleate at $x=L$, the BCs are adjusted to reflect this scenario. Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) posits that the void interface possesses an effective thickness, $\delta$, which is negligible relative to other dimensions within the interconnect tree. The examination of stress following void formation counts time, $t$, starting from the nucleation time, $t_{nuc}$. It is assumed that the initial stress distribution at this stage corresponds to the stress levels at the moment of void nucleation. This methodology facilitates the simulation of stress evolution within a single confined wire subsequent to the occurrence of void nucleation.
@@
$BC:\frac{\partial\sigma(0,t)}{\partial x}\bigg|_{x=0}=\frac{\sigma(0,t)}{\delta},\ t>0$
$BC:\frac{\partial\sigma(0,t)}{\partial x}\bigg|_{x=L}=-G,\ t>0$
@@
For analyzing multi-segment interconnects during the post-voiding phase, Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) is applied to the segment that contains the void. Segments without voids are addressed using Eq. (1). To accurately determine the stress profile across the entire interconnect tree, BCs that ensure stress continuity and atomic flux conservation must be implemented [27].
II-B Estimation of Variations in EM Stress Evolution
Conventional deterministic approach solves the combinations of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) to determine the post-voiding phase transient stress distribution, which averages the DC input current, thus does not account for variations in EM stress caused by fluctuations in input current, manufacturing processes, temperature, and diffusivity. Consequently, this method tends to yield overly optimistic and unrealistic predictions of component lifetimes [28].
If we define $\mathbf{x}$ as the vector of electrical and physical parameters influencing EM stress development, and $v$ as the error from these variations, the uncertainty and noise in these parameters can be treated as random variables within their respective spaces, i.e., $\mathbf{x}\in\mathbb{X}$ and $v\in\mathbb{V}$. It is posited that the solutions for EM stress are Gaussian distributed, centered around the true, but obscured, values [28] as shown in Eq. (3),